OpenStudy (elisaneedshelp):

Can I please be helped I really dont get this problem ):) Medal and Fan! Which expression is equivalent to this in its lowest terms?

OpenStudy (elisaneedshelp):

\[\sqrt{98x ^{8}y ^{5}z ^{40}}\]

OpenStudy (anonymous):

what is half of 8?

OpenStudy (elisaneedshelp):

4

OpenStudy (anonymous):

so an \(x^4\) comes right out front what is half of 40?

OpenStudy (elisaneedshelp):

20

OpenStudy (anonymous):

out comes a \(z^{20}\) how many times does 2 go in to 5?

OpenStudy (elisaneedshelp):

2 times

OpenStudy (anonymous):

with a remainder of?

OpenStudy (elisaneedshelp):

1

OpenStudy (anonymous):

ok good, so out comes \(y^2\) in stays \(y^1\) which is just \(y\)

OpenStudy (anonymous):

so far we are at \[x^4y^2z^{20}\sqrt{98y}\] now we take care of the \(98\)

OpenStudy (anonymous):

\[98=49\times 2\] so \[\sqrt{98}=\sqrt{49\times 2}=\sqrt{49}\sqrt{2}=7\sqrt2\]

OpenStudy (anonymous):

in other words out comes a 7, in stays a 2

OpenStudy (anonymous):

final answer \[7x^4y^2z^{20}\sqrt{2y}\] not too hard

OpenStudy (elisaneedshelp):

Thank you :) I did learn *thumbs up* !